OpenStudy (anonymous):

Can somebody please check if my work is correct for the following question: integral (2(xcosx-sinx))/x^2 =2 integral (xcosx-sinx)/x^2 =2 integral (xcosx)/x^2 - (sinx)/x^2 =2 integral cosx/x - 2 integral sinx/x^2 could somebody please explain it to me step by step?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

what is the next step if you could please show me

OpenStudy (anonymous):

do you know how to take integral by part?

OpenStudy (anonymous):

maybe i might remember if you showed me how

OpenStudy (anonymous):

but please explain it to me step by step

OpenStudy (anonymous):

for the first term, let u =1/x dv = cos x dx

OpenStudy (anonymous):

ok, let me do it on paper, you can see it clearer

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

That's it

OpenStudy (anonymous):

that is much better :)

OpenStudy (anonymous):

but i have a question

OpenStudy (anonymous):

You got it?

OpenStudy (anonymous):

cant we let u = cosx?

OpenStudy (anonymous):

yes i understood it thank you

OpenStudy (anonymous):

as i said cant we also solve it the other way around by letting u=cosx?

OpenStudy (anonymous):

You go nowhere on that way, because if you let u = cos x , du = -sin xdx dv = 1/x dx , v = lnx so that the second term is lnx cos x - \(\int \dfrac{lnx}{x}dx\) and then??? have to use substitute again for this second term. Moreover, it is useless for \(\int \dfrac{sinx}{x^2}dx\) You torture your life by going that way, hihihihihi

OpenStudy (anonymous):

hehehehehe

OpenStudy (anonymous):

but out of curiosity is it still right despite the torture lol

OpenStudy (anonymous):

ok, show me how to take \(\int \dfrac{sin x}{x^2}dx\)

OpenStudy (anonymous):

i would put it as integral 1/x^2 *sinx

OpenStudy (anonymous):

then integral x^-2 * sinx

OpenStudy (anonymous):

not to forget the dx at the end

OpenStudy (anonymous):

hihihihi... I save your time by posting the answer from mathtool website: http://www.wolframalpha.com/input/?i=int+%28sinx%2Fx^2%29+dx

OpenStudy (anonymous):

what does Ci(x) stand for?

OpenStudy (anonymous):

I don't know, it's out of our level. It may be for graduate level.

OpenStudy (anonymous):

but even when i tried to integrate x^-2 * sinx i got x^-1/-1 * (-cosx)

OpenStudy (anonymous):

which equals to cosx/x

OpenStudy (anonymous):

i dont know where i got it wrong

OpenStudy (anonymous):

nope, you let u = x^(-2) , right? then du = -2x^(-3) hehehe.... the integral becomes bigger and bigger.

OpenStudy (anonymous):

ok so that would be my u

OpenStudy (anonymous):

if i had to solve it

OpenStudy (anonymous):

is there a trick to know which is the u?

OpenStudy (anonymous):

ill let u know when i get back. thanks for your help :)

OpenStudy (anonymous):

i now understand why thanks to you :) I really appreciate your help

OpenStudy (anonymous):

can it be written like this: 2 integral cosx/x dx -2 integral sinx/x^2 dx 2 integral (sinx/x + integral sinx/x^2) -2 integral sinx/x^2 dx

OpenStudy (anonymous):

@OOOPS

OpenStudy (anonymous):

could ypu please help me @OOOPS